# BlackboxBench: A Comprehensive Benchmark of Black-box Adversarial Attacks

### v1

@@ IV-A Experimental Setup
Models In query-based attacks, we attack 4(5) normally trained target models and 2(3) adversarially trained target models on CIFAR-10(ImageNet) dataset. In transfer-based attacks, following prior works[30, 39, 35, 29, 42], we choose 4(5) surrogate models, evaluate 8(14) normally trained target models and 1(3) adversarially trained target models on CIFAR-10(ImageNet) dataset. The model architectures are diverse and belong to different families to study the transferability. All the adopted models and their clean accuracies on benign test sets are shown in Tab. III. On CIFAR-10, they are evaluated on 10,000 test set. On ImageNet, they are evaluated on the 50,000 validation set. The download links of checkpoints are provided in Appendix A-A.

### v2

@@ IV-A Experimental Setup
Models In query-based attacks, we attack 4 normally trained target models and 2 adversarially trained target models on the CIFAR-10 dataset, and 5 normally trained target models and 3 adversarially trained target models on the ImageNet dataset. In transfer-based attacks, following prior works [66, 4, 72, 65, 16], we choose 4 surrogate models, evaluate 8 normally trained target models and 1 adversarially trained target model on the CIFAR-10 dataset, and choose 5 surrogate models, evaluate 14 normally trained target models and 3 adversarially trained target models on the ImageNet dataset. The model architectures are diverse and belong to different families to study the transferability. The clean accuracies and checkpoints of adopted models are provided in the Supplementary D-A.
@@ IV-B Result Overview
[…]
In query-based attacks, we evaluate 29 methods outlined in Supplementary A using datasets and models detailed in Sec. IV-A across all four attack settings. $\ell_{2}$ and $\ell_{\infty}$ are implemented for most attacks, even though some attacks are explicitly designed for one certain norm. There are a total of 630 pairs of evaluations for decision-based black-box attacks and 988 pairs of evaluations for score-based black-box attacks. Each method’s performance in attacking a target model is assessed using three metrics: average query number (AQN), median query number (MQN), and attack success rate (ASR). A better attacker should report a higher ASR and a lower AQN and MQN simultaneously. In transfer-based attacks, we evaluate 30 methods mentioned in Supplementary A on datasets and models listed in Sec. IV-A under all four attack settings, leading to 13,332 evaluations in total. The efficacy of adversarial perturbations crafted on a surrogate model by each method is measured by averaging ASR on all targeted models. A superior attack method is expected to yield a higher averaged ASR. For the context of transfer-based attacks, unless explicitly specified otherwise, "ASR" refers to the averaged ASR herein.
Due to space limitations, in Fig. 5 and Fig. 6, we exclusively offer a condensed summary of results specifically for the ImageNet dataset alongside one model architecture—ResNet-50, $i.e.$, we only exhibit results where ResNet50 acts as a surrogate model in transfer-based attacks and as target model in query-based attacks. For comprehensive tables including additional models and the CIFAR-10 dataset, please refer to BlackboxBench LeaderBoard. Based on the results in Fig. 5 and Fig. 6, for each type of attack, our overview in this section will be organized from the following perspectives: 1) what is the progress in the field over the past years? 2) which category presents the most promising avenue? 3) under each attack setting, what is the best-performing method in each category and what is the SOTA method? 4) can individual attacks be combined to form a stronger attack? If so, what is the most effective guideline for combination?
Note: For question 4), we will omit discussions on decision-based and score-based attacks, focusing solely on the case of transfer-based attacks for following reasons. Since decision-based and score-based attacks rely on limited information from the target model (only hard labels or confidence scores) and are constrained by query limits, existing works mainly focus on designing the search strategy to achieve effective attacks with minimal queries. In contrast, transfer-based attacks leverage full information from surrogate models (such as gradients) and the advantage of being unrestricted by query limits, making it possible to combine multiple strategies to significantly enhance attack success rates and transferability.
IV-B1 Overview of decision-based attacks